OpenStudy (anonymous):

Raymond is buying boxes of party favors for his son birthday party. he has to decide how many boxes he should purchase. A formula for this scenario is: c=nz/f c=number of children at the party n=number of party favors per box z=number of boxes f=number of favors each child will get Rewrite the equation to solve for the number of boxes of party favors Raymond should purchase.

OpenStudy (anonymous):

Hi @mindofastoner_drea What we need to do here is rearrange the equation. By looking at the variables, we can see that we need to isolate for z, the total number of boxes. First, let's multiply both sides by f ... c*f = n*z Then, let's divide both sides by n in order to isolate z on the right hand side... c*f/n = z. Let me know if you don't understand any of these steps!

OpenStudy (anonymous):

Thank you can you also help me with 5.5=m/n I have to isolate the m. How do i do that? The same way?

OpenStudy (anonymous):

@Nengeni

OpenStudy (anonymous):

@mindofastoner_drea What you need to do is get the m all alone on one side of the equation. Do you see anything that you can do to the right hand side that will leave you with only an m?

OpenStudy (anonymous):

multiply?

OpenStudy (anonymous):

Sorry! I didn't get a notification! You are absolutely correct! If you multiply both sides by n, you will end up with 5.5n = m!